# Property prediction for ionic liquids without prior structural knowledge using limited experimental data: A data-driven neural recommender system leveraging transfer learning

### v1

@@ Abstract
Ionic liquids (ILs) have emerged as versatile replacements for traditional solvents because their physicochemical properties can be precisely tailored to various applications. However, accurately predicting key thermophysical properties remains challenging due to the vast chemical design space and the limited availability of experimental data. In this study, we present a data-driven transfer learning framework that leverages a neural recommender system (NRS) to enable reliable property prediction for ILs using sparse experimental datasets. The approach involves a two-stage process: first, pre-training NRS models on COSMO-RS-based simulated data at fixed temperature and pressure to learn property-specific structural embeddings for cations and anions; and second, fine-tuning simple feedforward neural networks using these embeddings with experimental data at varying temperatures and pressures. In this work, five essential IL properties are considered: density, viscosity, surface tension, heat capacity, and melting point. The framework supports both within-property and cross-property knowledge transfer. Notably, pre-trained models for density, viscosity, and heat capacity are used to fine-tune models for all five target properties, achieving improved performance by a substantial margin for four of them. The model exhibits robust extrapolation to previously unseen ILs. Moreover, the final trained models enable property prediction for over 700,000 IL combinations, offering a scalable solution for IL screening in process design. This work highlights the effectiveness of combining simulated data and transfer learning to overcome sparsity in the experimental data.
@@
[…]
[a]organization=Otto-von-Guericke University Magdeburg, Chair of Process Systems Engineering, addressline=Universitätsplatz 2, city=Magdeburg, postcode=39106, country=Germany
[…]
@@ 1 Introduction
Ionic liquids (ILs) are a class of salts that remain liquid at or near room temperature and have gained attention as potential alternatives to conventional solvents (Brennecke and Maginn, 2001; Welton, 1999). Their favorable properties, including low vapor pressure, high thermal stability, and a wide liquid temperature range, make them suitable for a variety of chemical engineering applications (Patel and Lee, 2012). In addition, ILs are structurally versatile and their physicochemical properties can be adjusted by selecting different combinations of cations and anions. This tunability offers the opportunity to design ILs tailored to specific process requirements. However, practical use of ILs requires knowledge of key properties such as density, viscosity, surface tension, heat capacity, and melting point. Experimentally determining these properties for the vast number of potential ILs is expensive and time-consuming. Therefore, predictive modeling has become an essential tool for screening ILs and guiding their selection in process design.
@@
In the context of designing and selecting ILs for chemical processes, several fundamental physical properties play a critical role in determining their practical applicability and process performance. Density is essential for material and energy balances, equipment design, and phase behavior prediction (Birdi, 1997). Viscosity influences the flow behavior of ILs, affecting mass and heat transfer, mixing, and pumping requirements (Shen and Nicolaou, 2019). Surface tension governs interfacial interactions and is especially relevant in applications involving liquid–liquid extraction, wetting, or emulsification (Baskin et al., 2022a). Heat capacity reflects a fluid’s ability to store thermal energy and is essential for thermal management, energy integration, and accurate energy balance calculations in process modeling. Finally, melting point sets the lower operational limit for ILs and is used to exclude candidates that may solidify under working conditions. Reliable knowledge or prediction of these properties is vital for effective process modeling, simulation, optimization, and making informed process-related decisions. However, due to the enormous variety of possible IL structures, identifying suitable candidates with the right combination of properties remains a complex and computationally intensive task (Wasserscheid and Keim, 2000; Stark and Seddon, 2000).
@@
Traditionally, the identification of suitable ionic liquids for process applications has relied heavily on experimental screening, which is often slow, costly, and restricted to a narrow set of known IL structures. As a result, a large portion of the IL design space remains unexplored. To overcome these limitations, computational approaches have been increasingly used to estimate key thermophysical properties of ILs (Sethi et al., 2025; Sengers et al., 2000). Classical methods such as equations of state (EoS) (Turner et al., 2003) and group contribution methods (GCMs) (Boudouris et al., 1997) have been widely applied to predict properties like density, viscosity, and heat capacity. However, these approaches face important limitations. Equations of state often require parameter fitting and may not generalize well to new IL structures, while GCMs depend on the availability and quality of group parameters, which are sometimes missing or inaccurate for novel ionic liquids. As a result, both techniques may lack the flexibility and precision needed to predict a broad set of IL properties across a wide structural range.
@@
One of the widely adopted strategies for predicting the properties of ILs is the quantitative structure–property relationship (QSPR) approach. This method builds statistical models that associate target properties with structural features of ILs, typically encoded through molecular descriptors (Sepehri, 2020; Rybinska-Fryca et al., 2020). With the growth of publicly available IL property databases such as PubChem (Kim et al., 2025) and ILThermo (Dong et al., 2007), a wide variety of descriptors ranging from simple topological indices to complex quantum chemical parameters have been used to represent ILs. In recent years, machine learning techniques have increasingly been employed within QSPR frameworks, enhancing their flexibility and predictive power for modeling diverse IL properties (Cao et al., 2018; Chen et al., 2024; Sun et al., 2021; Ding et al., 2021; Wu et al., 2024; Baskin et al., 2022b; Mohan et al., 2024). However, these descriptor-based models still depend heavily on the quality and completeness of the descriptors themselves, which are often designed manually and tailored to specific tasks (Han et al., 2011; Rybinska et al., 2016; Xia et al., 2023; Lotfi et al., 2021). This limits their ability to generalize across structurally diverse IL families. Graph neural networks (GNNs) have been explored for IL property prediction, leveraging molecular graph representations to capture structural relationships with no reliance on traditional molecular descriptors (Sanchez Medina and Sundmacher, 2026; Rittig et al., 2023; Qin et al., 2023). However, GNNs often require large datasets to effectively learn these representations, limiting their performance when experimental data is scarce. As an alternative, matrix completion methods based on neural networks have emerged (Chen et al., 2021; Zhang et al., 2022; Ramlatchan et al., 2018; Athey et al., 2021; Fan and Cheng, 2018; Fan and Chow, 2017). These methods treat the IL property dataset as a matrix with missing entries and aim to infer the unknown values by exploiting correlations between ILs and between properties. This approach eliminates the need for explicit descriptors and instead learns latent patterns directly from available data. However, such models require a sufficiently large amount of labeled data to train effectively and to generalize to unseen ILs (Chen et al., 2021; Zhang et al., 2022). In cases of limited experimental data, they often struggle with overfitting or fail to capture meaningful structure-property relationships.
@@
To address the challenge of data sparsity in property prediction, transfer learning has recently emerged as a promising solution (Weiss et al., 2016; Pan and Yang, 2010; Dai et al., 2007; Neyshabur et al., 2020). This technique has so far been utilized by pre-training models on large-scale unlabeled data to learn general molecular representations, which are then fine-tuned on relatively small labeled datasets for specific prediction tasks. In the context of ionic liquids (ILs), one recent study has explored this approach by leveraging extensive unlabeled SMILES data to derive structural embeddings of IL ions, bypassing the need for manually crafted molecular descriptors (Chen et al., 2023). These learned representations were subsequently fine-tuned using limited experimental data to predict various thermophysical properties. While this framework has demonstrated improved generalization and superior performance compared to conventional models, its accuracy notably declines in the presence of extremely sparse experimental data, such as for heat capacity and surface tension. This limitation highlights the need for more robust strategies to ensure reliable predictions across a broader spectrum of IL properties.
@@
In this work, we propose a transfer learning framework that leverages a neural recommender system (NRS) to overcome the limitations posed by sparse experimental data in predicting key physicochemical properties of ILs. The NRS is employed during the pre-training phase to learn property-specific structural embeddings from COSMO-based simulation data for density, viscosity, and heat capacity at fixed temperature and pressure. These embeddings capture essential molecular features without the confounding influence of thermodynamic variables. In the subsequent fine-tuning phase, a simple feedforward neural network is trained for each target property (density, viscosity, surface tension, heat capacity, and melting point) using the learned structural embeddings as input and experimental property data for supervision. This two-stage framework allows both within-property and cross-property knowledge transfer, enabling embeddings learned from one property to improve predictions for others. The results demonstrate that this strategy achieves high predictive accuracy even for properties with extremely limited experimental data, confirming the robustness and transferability of the learned representations.
@@
This study is organized as follows: In section 2, we detail the methodology, including the neural recommender system and transfer learning framework, the pre-training phase using COSMO-RS simulated data, and the fine-tuning phase with experimental data. Section 3 presents the datasets, covering the generation of COSMO-based molecular structures, data sampling, and the collection of experimental data from literature. Section 4 discusses the results, evaluating model performance and the impact of the size of the pre-training dataset on prediction accuracy. Finally, section 4 summarizes the conclusions and potential for future application of the proposed methodology.
@@ 2.1 Neural recommender system and transfer learning framework
In this work, we implement a two-step transfer learning framework. An overview of the complete workflow is illustrated in LABEL:fig:workflow. Separate neural recommender system models are initially pre-trained for each of the three properties: density, viscosity, and heat capacity. This pre-training is conducted using COSMO-generated simulated data at fixed temperature and pressure ($298\text{\,}\mathrm{K}$ and $1\text{\,}\mathrm{bar}$). The goal is to learn property-specific, meaningful structural embeddings for each cation and anion by capturing their interactions and extracting latent features that represent a reduced molecular structural space conditioned on the respective property. These learned embeddings are subsequently fixed and reused, while a feed-forward neural network composed exclusively of dense layers is fine-tuned using sparse experimental data. This setup focuses on modeling the influence of the structural representations on each property under varying temperature and pressure. By sequentially learning structural embeddings and the effects of temperature and pressure, this approach effectively overcomes major challenges in IL property prediction, particularly addressing the limited and imbalanced experimental dataset.
@@ 2.1.1 Pre-training phase
In the pre-training phase, a neural recommender system is developed to learn property-specific structural embeddings for ILs, as shown in LABEL:fig:nrs_architecture. Cations and anions are provided as index-based inputs, which are mapped to the target property at fixed temperature and pressure ($298\text{\,}\mathrm{K}$ and $1\text{\,}\mathrm{bar}$) to learn latent vectors of dimension 100, representing structural embeddings for each ion. The embedding vectors for cation and anion are independently processed through a dense neural block resembling a residual network (ResNet) architecture (Targ et al., 2016). Each block consists of a series of dense layers with varying number of neurons selected from [100, 200, 300], followed by dropout (rate = 0.05) and element-wise addition layers. The outputs of the individual cation and anion pathways are then concatenated to form a combined latent representation. This concatenated vector is further passed through two additional fully connected layers, the first with a neuron size selected from [50, 100, 200, 400], and the second with half the number of neurons of the preceding layer. Rectified Linear Unit (ReLU) activation functions are applied to all dense layers, while the output layer remains activation-free to prevent range constraints on predicted values. The model is trained using the Adam optimizer with a learning rate of 0.001.
Pre-training is carried out separately for three target properties - density, viscosity, and heat capacity - using identical model architectures and optimization strategies for each property. Hyperparameters related to the number of neurons are tuned individually per property. Pre-training is not performed for surface tension as no corresponding COSMO-RS model is available. Melting point is also excluded from the pre-training phase since the COSMO-RS based simulation dataset contains numerous missing values. These gaps primarily arise from convergence issues during the quantum chemical optimization of certain cation–anion structures, which are required for generating reliable COSMO descriptors. To avoid data leakage, it is ensured that ILs included in any of the experimental datasets for the five properties are excluded from all pre-training datasets. Furthermore, rigorous ten-fold cross-validation is conducted during pre-training to select optimal hyperparameters. In each fold, train-test splits are performed rigorously based on ILs rather than randomly on individual data points, ensuring that no ionic liquid appears in both training and test sets.
@@ 2.1.2 Fine-tuning phase
In the fine-tuning phase, the pre-trained components of the neural recommender system corresponding to cation and anion representations, up to and including the concatenation layer, are reused with fixed weights and treated as a non-trainable forward pass, as shown in LABEL:fig:nrs_architecture. These embeddings remain constant during fine-tuning. Additionally, temperature and/or pressure inputs are concatenated with the fixed ion embeddings to form the final input representation. This combined vector is then passed through two fully trainable dense layers. The first layer uses a neuron size selected from [50, 100, 200], while the second layer contains half the number of neurons as the preceding layer. A dropout rate of 0.05 is applied after each dense layer to mitigate overfitting. The Adam optimizer is employed for training with a learning rate of 0.01.
@@
The structural embeddings learned from pre-trained models for the three properties, namely density, viscosity, and heat capacity, are each used as a basis for fine-tuning all five target properties, including surface tension and melting point. This approach enables the evaluation of both within-property and cross-property transfer learning. The performance outcomes across different source–target property combinations are systematically compared.
As illustrated in Figure 3, the paucity of experimental data highlights the challenges in property prediction and emphasizes the necessity of the transfer learning approach. This methodology enables extrapolation across the entire chemical space of 2,268 cations and 311 anions included in the study. By learning structural embeddings from unbiased simulated data during pre-training, the fine-tuned model not only extrapolates to ILs unseen during training but also predicts properties for combinations where both the cation and the anion are absent from the experimental data.
@@ 2.2.1 Generation of molecular structures via DFT calculations
First, all unique cations and anions present in the experimental datasets for the five target properties (as listed in LABEL:tab:performance_comparison_1) are identified. Canonical SMILES strings for each ion are generated using RDKit (Landrum, 2013). These SMILES strings are then used to create optimized molecular geometries and COSMO files via a series of quantum chemical calculations. All Density Functional Theory (DFT) calculations are carried out using TURBOMOLE, version 7.7.1 (Furche et al., 2014). Chemical structures are optimized with the BP86 GGA (generalized gradient approximation) functional (Becke, 1988; Perdew, 1986), including DFT-D3 (Grimme et al., 2010), dispersion corrections with Becke-Johnson (BJ) damping (Grimme et al., 2011), and using Ahlrichs’ triple-$\zeta$ valence polarization basis set (def2-TZVP) (Schäfer et al., 1994) for all atoms.
The input files for COSMO-RS (conductor like screening model for real solvents) (Klamt, 1995) are generated via single point calculations at BP86-D3(BJ)/def2-TZVPD level of theory, setting an infinite permittivity and using the refined COSMO cavity construction algorithm. The resulting files form the structural input required for property prediction via COSMO-RS.

### v2

# A neural recommender system leveraging transfer learning for property prediction of ionic liquids

@@ Abstract
Ionic liquids (ILs) have emerged as versatile replacements for traditional solvents because their physicochemical properties can be precisely tailored to various applications. However, accurately predicting key thermophysical properties remains challenging due to the vast chemical design space and the limited availability of experimental data. In this study, we present a data-driven transfer learning framework combined with a neural recommender system (NRS) to enable reliable property prediction for ILs using sparse experimental datasets. The approach involves a two-stage process: first, pre-training NRS models on COSMO-RS-based simulated data at fixed temperature and pressure, and second, fine-tuning simple feedforward neural networks with experimental data at varying temperatures and pressures. In this work, five essential IL properties are considered: density, viscosity, surface tension, heat capacity, and melting point. We find that the framework supports both within-property and cross-property knowledge transfer. Notably, pre-trained models for density, viscosity, and heat capacity are used to fine-tune models for all five target properties, achieving improved performance by a substantial margin for four of them. The model exhibits robust extrapolation to previously unseen ILs. Moreover, the final trained models enable property prediction for over 700,000 IL combinations, offering a scalable solution for IL screening in process design. This work highlights the effectiveness of combining simulated data and transfer learning to overcome sparsity in the experimental data.
@@
[…]
[a]organization=Otto von Guericke University Magdeburg, Chair of Process Systems Engineering, addressline=Universitätsplatz 2, city=Magdeburg, postcode=39106, country=Germany
[…]
@@ 1 Introduction
Ionic liquids (ILs) are a class of salts that remain in the liquid state at or near room temperature and have gained attention as potential alternatives to conventional solvents (Brennecke2001; Welton1999). Their favorable properties, including low vapor pressure, high thermal stability, and a wide liquid temperature range, make them suitable for a variety of process engineering applications (Patel2012). In addition, ILs are structurally versatile and their physicochemical properties can be adjusted by selecting different combinations of cations and anions. This tunability offers the opportunity to design ILs tailored to specific process requirements. However, practical use of ILs requires knowledge of key properties such as density, viscosity, surface tension, heat capacity, and melting point. Experimentally determining these properties for the vast number of potential ILs is expensive and time-consuming. Therefore, predictive modeling has become an essential tool for screening ILs and guiding their selection in process design.
@@
In the context of designing and selecting ILs for chemical processes, several fundamental physical properties play a critical role in determining their practical applicability and process performance. Density is essential for material and energy balances, equipment design, and phase behavior prediction (Birdi2016). Viscosity influences the flow behavior of ILs, affecting mass and heat transfer, mixing, and pumping requirements (Shen2019). Surface tension governs interfacial interactions and is especially relevant in applications involving liquid–liquid extraction, wetting, or emulsification (Baskin2022). Heat capacity reflects a fluid’s ability to store thermal energy and is essential for thermal management, energy integration, and accurate energy balance calculations in process modeling. Finally, melting point sets the lower operational limit for ILs and is used to exclude candidates that may solidify under working conditions. Reliable knowledge or prediction of these properties is vital for effective process modeling, simulation, optimization, and making informed process-related decisions. However, due to the enormous variety of possible IL structures, identifying suitable candidates with the right combination of properties remains a complex and computationally intensive task (Wasserscheid2000; Stark2000).
@@
Traditionally, the identification of suitable ionic liquids for process applications has relied heavily on experimental screening, which is often slow, costly, and restricted to a narrow set of known IL structures. As a result, a large portion of the IL design space remains unexplored. To overcome these limitations, computational approaches have been increasingly used to estimate key thermophysical properties of ILs (Sethi2025; Sengers2000). Classical methods such as equations of state (EoS) (Turner2003) and group contribution methods (GCMs) (Gani1983) have been widely applied to predict properties like density, viscosity, and heat capacity. However, these approaches face important limitations. Equations of state often require parameter fitting and may not be generalizable to new IL structures, while GCMs depend on the availability and quality of group parameters, which are sometimes missing or inaccurate for novel ionic liquids. As a result, both techniques may lack the flexibility and precision needed to predict a broad set of IL properties across a wide structural range.
@@
One of the widely adopted strategies for predicting the properties of ILs is the quantitative structure–property relationship (QSPR) approach. This method builds statistical models that associate target properties with structural features of ILs, typically encoded through molecular descriptors (Sepehri2020; RybinskaFryca2020). With the growth of publicly available IL property databases such as PubChem (kim2025) and ILThermo (dong2007), a wide variety of descriptors ranging from simple topological indices to complex quantum chemical parameters have been used to represent ILs. In recent years, machine learning techniques have increasingly been employed within QSPR frameworks, enhancing their flexibility and predictive power for modeling diverse IL properties (Cao2018; Chen2024; Sun2021; Ding2021; Wu2024; Baskin2022b; Mohan2024). However, these descriptor-based models still depend heavily on the quality and completeness of the descriptors themselves, which are often designed manually and tailored to specific tasks (Han2011; Rybinska2016; Xia2023; Lotfi2021). This limits their ability to generalize across structurally diverse IL families. Graph neural networks (GNNs) have been explored for IL property prediction, leveraging molecular graph representations to capture structural relationships with no reliance on traditional molecular descriptors (sanchez2026graph; rittig2023graph; qin2023capturing). However, GNNs often require large datasets to effectively learn these representations, limiting their applicability when experimental data is scarce. As an alternative, matrix completion methods based on neural networks have emerged (Chen2021; Zhang2022; Ramlatchan2018; Athey2021; Fan2018; Fan2017). These methods treat the IL property dataset as a matrix with missing entries and aim to infer the unknown values by exploiting correlations between ILs and between properties. This approach eliminates the need for explicit descriptors and instead learns latent patterns directly from available data. However, such models require a sufficiently large amount of labeled data to train effectively and to predict the properties of ILs unseen in the training step. (Chen2021; Zhang2022). In cases of limited experimental data, they often struggle with overfitting or fail to capture meaningful structure-property relationships.
@@
To address the challenge of data sparsity in property prediction, transfer learning has recently emerged as a promising solution (Weiss2016; Dai2007; Neyshabur2020). This technique has so far been utilized by pre-training models on large-scale unlabeled data to learn general molecular representations, which are then fine-tuned on relatively small labeled datasets for specific prediction tasks. In the context of ionic liquids (ILs), one recent study has explored this approach by leveraging extensive unlabeled SMILES data to derive structural embeddings of IL ions, bypassing the need for manually crafted molecular descriptors (Chen2023). These learned representations were subsequently fine-tuned using limited experimental data to predict various thermophysical properties. While this framework has demonstrated improved generalization and superior performance compared to conventional models, its accuracy notably declines in the presence of extremely sparse experimental data, such as for heat capacity and surface tension. This limitation highlights the need for more robust strategies to ensure reliable predictions across a broader spectrum of IL properties.
@@
In this work, we propose a transfer learning framework that leverages a neural recommender system (NRS) to overcome the limitations posed by sparse experimental data in predicting key physicochemical properties of ILs. The NRS is employed during the pre-training phase to learn property-specific structural embeddings from COSMO-based simulation data for density, viscosity, and heat capacity at fixed temperature and pressure. These embeddings capture essential molecular features without the confounding influence of thermodynamic variables. In the subsequent fine-tuning phase, a simple feedforward neural network is trained for each target property (density, viscosity, surface tension, heat capacity, and melting point) using the learned structural embeddings as inputs and experimental property data for supervision.
@@
This study is organized as follows: In section 2, we detail the methodology, including the neural recommender system and transfer learning framework, the pre-training phase using COSMO-RS simulated data, and the fine-tuning phase with experimental data. This section also presents the datasets, covering the generation of COSMO-based molecular structures, data sampling, and the collection of experimental data from literature. Section 3 discusses the results, evaluating model performance and the impact of the size of the pre-training dataset on prediction accuracy. Finally, section 4 summarizes the conclusions and potential for future application of the proposed methodology.
@@ 2.1 Neural recommender system and transfer learning framework
In this work, we implement a two-step transfer learning framework. An overview of the complete workflow is illustrated in LABEL:fig:workflow. Separate neural recommender system models are initially pre-trained for each of the three properties: density, viscosity, and heat capacity. This pre-training is conducted using COSMO-generated simulated data at fixed temperature and pressure ($298\text{\,}\mathrm{K}$ and $1\text{\,}\mathrm{bar}$). The goal is to learn property-specific, meaningful structural embeddings for each cation and anion by capturing their interactions and extracting latent features that represent a reduced molecular structural space conditioned on the respective property. These learned embeddings are subsequently fixed and reused, while a feed-forward neural network composed exclusively of dense layers is fine-tuned using sparse experimental data. This setup focuses on modeling the influence of the structural representations on each property under varying temperature and pressure. By sequentially learning structural embeddings and the effects of temperature and pressure, this approach can effectively overcome major challenges in IL property prediction, particularly addressing the limited and imbalanced experimental dataset.
@@ 2.1.1 Pre-training phase
In the pre-training phase, a neural recommender system is developed to learn property-specific structural embeddings for ILs, as shown in LABEL:fig:nrs_architecture. Cations and anions are provided as index-based inputs, which are mapped to the target property at fixed temperature and pressure ($298\text{\,}\mathrm{K}$ and $1\text{\,}\mathrm{bar}$) to learn latent vectors of dimension 100, representing structural embeddings for each ion. The embedding vectors for cation and anion are independently processed through a dense neural block resembling a residual network (ResNet) architecture (targ2016resnet). Each block consists of a series of dense layers with varying number of neurons selected from [100, 200, 300], followed by dropout (rate = 0.05) and element-wise addition layers. The outputs of the individual cation and anion pathways are then concatenated to form a combined latent representation. The concatenated vector is subsequently passed through two additional fully connected layers, as one layer is found to be insufficient, the first with a number of neurons selected from [50, 100, 200, 400], and the second with half the number of neurons of the preceding layer. Rectified Linear Unit (ReLU) activation functions are applied to all dense layers, while the output layer remains activation-free to prevent range constraints on predicted values. The model is trained using the Adam optimizer (kingma2014adam) with a learning rate of 0.001.
Pre-training is carried out separately for three target properties - density, viscosity, and heat capacity - using identical model architectures and optimization strategies for each property. Hyperparameters related to the number of neurons are tuned individually per property. Pre-training is not performed for surface tension as no corresponding COSMO-RS model is available. Melting point data is also excluded from the pre-training phase since numerous values are missing in the COSMO-RS-based simulation dataset. These gaps primarily arise from convergence issues during the quantum chemical optimization of certain cation–anion structures, which are required for generating reliable COSMO descriptors. To avoid data leakage, it is ensured that ILs included in any of the experimental datasets for the five properties are excluded from all pre-training datasets. Furthermore, rigorous ten-fold cross-validation is conducted during pre-training to select optimal hyperparameters. In each fold, train-test splits are performed rigorously based on ILs rather than randomly on individual data points, ensuring that no ionic liquid appears in both training and test sets.
@@ 2.1.2 Fine-tuning phase
In the fine-tuning phase, the pre-trained components of the neural recommender system corresponding to cation and anion representations, up to and including the concatenation layer, are reused with fixed weights and treated as a non-trainable forward pass, as shown in LABEL:fig:nrs_architecture. These embeddings remain constant during fine-tuning. Additionally, temperature and/or pressure inputs are concatenated with the fixed ion embeddings to form the final input representation. This combined vector is then passed through two fully trainable dense layers. The first layer uses a number of neurons selected from [50, 100, 200], while the second layer contains half the number of neurons as the preceding layer. A dropout rate of 0.05 is applied after each dense layer to mitigate overfitting. The Adam optimizer is employed for training with a learning rate of 0.01.
@@
The structural embeddings learned from pre-trained models for the three properties, namely density, viscosity, and heat capacity, are each used as a basis for fine-tuning of all five target properties, including surface tension and melting point. This approach enables the evaluation of both within-property and cross-property transfer learning. The performance outcomes across different source–target property combinations are systematically compared.
As illustrated in Figure 3, the paucity of experimental data highlights the challenges in property prediction and emphasizes the necessity of the transfer learning approach. This methodology enables extrapolation across the entire chemical space of 2,268 cations and 311 anions included in the study. By learning structural embeddings from unbiased simulated data during pre-training, the fine-tuned model not only extrapolates to ILs unseen during training but also predicts properties for combinations where both the cation and the anion are not contained in from the experimental data.
@@ 2.2.1 Generation of molecular structures via DFT calculations
First, all unique cations and anions present in the experimental datasets for the five target properties (as listed in LABEL:tab:performance_comparison_1) are identified. Canonical SMILES strings for each ion are generated using RDKit (landrum2013rdkit). These SMILES strings are then used to create optimized molecular geometries and COSMO files via a series of quantum chemical calculations. All Density Functional Theory (DFT) calculations are carried out using TURBOMOLE, version 7.7.1 (Furche2014). Chemical structures are optimized with the BP86 GGA (generalized gradient approximation) functional (Becke1988; Perdew1986), including DFT-D3 (Grimme2010), dispersion corrections with Becke-Johnson (BJ) damping (Grimme2011), and using Ahlrichs’ triple-$\zeta$ valence polarization basis set (def2-TZVP) (Schafer1994) for all atoms.
The input files for COSMO-RS (conductor like screening model for real solvents) (Klamt1995) are generated via single point calculations at BP86-D3(BJ)/def2-TZVPD level of theory, setting an infinite permittivity and using the refined COSMO cavity construction algorithm. The resulting files form the structural input required for property prediction via COSMO-RS.